OpenStudy (anonymous):

Solve the following system of equations. 3x+2y=-12 x=2

OpenStudy (anonymous):

You are given x so Plug x into the equation and solve for y.

OpenStudy (the_fizicx99):

Your are given the x value so plug it in 3(2)+2y=-12 6 + 2y = -12 Solve for y

OpenStudy (anonymous):

Some of my answers are (2,-9), (-2,-3),(3,2) and (2,3) and I will give a medal

OpenStudy (the_fizicx99):

Solve for y lol.

OpenStudy (the_fizicx99):

ISOLATE Y, subtract 6 from both sides then divide by 2

OpenStudy (anonymous):

Sill confused...

OpenStudy (the_fizicx99):

Ermagawd 6 + 2y = -12 -6 -6 2y = -18 /2 /2 y = ??

OpenStudy (anonymous):

is it 2,-9

OpenStudy (the_fizicx99):

ye